# All-weather Multi-Modality Image Fusion: Unified Framework and 100k Benchmark

### v2

@@ Abstract
Multi-modality image fusion (MMIF) combines complementary information from different image modalities to provide a more comprehensive and objective interpretation of scenes. However, existing MMIF methods lack the ability to resist different weather interferences in real-world scenes, preventing them from being useful in practical applications such as autonomous driving. To bridge this research gap, we proposed an all-weather MMIF model. Achieving effective multi-tasking in this context is particularly challenging due to the complex and diverse nature of weather conditions. A key obstacle lies in the 'black box' nature of current deep learning architectures, which restricts their multi-tasking capabilities. To overcome this, we decompose the network into two modules: a fusion module and a restoration module. For the fusion module, we introduce a learnable low-rank representation model to decompose images into low-rank and sparse components. This interpretable feature separation allows us to better observe and understand images. For the restoration module, we propose a physically-aware clear feature prediction module based on an atmospheric scattering model that can deduce variations in light transmittance from both scene illumination and reflectance. We also construct a large-scale multi-modality dataset with 100,000 image pairs across rain, haze, and snow conditions, covering various degradation levels and diverse scenes to thoroughly evaluate image fusion methods in adverse weather. Experimental results in both real-world and synthetic scenes show that the proposed algorithm excels in detail recovery and multi-modality feature extraction. The code is available at https://github.com/ixilai/AWFusion.
Index Terms:
@@
Image fusion, adverse weather, multi-modality benchmark, image restoration.I Introduction
@@
Infrared and visible image fusion (IVIF) is performed to integrate the complementary and salient feature information from different modal images to achieve a more comprehensive interpretation of a target scene. By integrating data from different spectral ranges, IVIF contributes to advancements in computer vision tasks [1, 2, 3, 4, 5], including semantic segmentation [6, 7, 8], object detection [9, 10, 11], and unmanned driving [12]. Recently, deep learning (DL) has exhibited remarkable progress within the realm of artificial intelligence owing to its exceptional aptitude for intricate nonlinear approximations and adeptness in extracting salient features. Consequently, several highly efficient algorithms tailored to the IVIF domain have been developed. These algorithms can be broadly categorized into three groups: generative adversarial networks (GAN)-based models [13, 14, 15, 16], autoencoder (AE)-based models [2, 17, 18], and algorithmic expansion models [4, 19].
Each of these methodologies has achieved commendable outcomes in the IVIF field. However, there are three key questions that remain unresolved.
@@
Currently, there is a critical research gap in that IVIF cannot deliver robust scene information under adverse weather conditions. Addressing this challenge requires more than isolated algorithmic solutions for specific weather-related issues, necessitating a unified framework to comprehensively restore intricate scene details affected by adverse weather.
@@
Generally, DL-based algorithms are driven by data. Existing datasets dedicated to collecting comprehensive imaging data from multiple modalities are complete in terms of scene diversity. However, no large datasets have been developed specifically for research on applying IVIF to scenarios with adverse weather conditions. Even if certain datasets can cover scenes affected by rain and haze, their limited data make it difficult to employ them for training in joint image restoration and fusion.
To address above challenges, we propose an all-weather MMIF algorithm (AWFusion) that integrates both deweathering and fusion functions within a single framework, thereby providing an effective means to achieve robust fusion under adverse weather conditions. Compared to the traditional two-stage strategy of ‘restoration+ fusion’ for addressing adverse weather fusion problems, we propose that an end-to-end framework, which prioritizes modality information interaction before feature recovery, can yield more effective results. The two-stage approach has notable drawbacks: because image restoration and fusion are independent processes, errors introduced during restoration directly impact the fusion outcome. Additionally, the lack of correlation between these tasks means that while efficient restoration may enhance fusion performance, the reverse is not true.
@@
In complex natural scenes, images often contain substantial redundant information, allowing for the inference of the original scene even when the image quality is degraded due to adverse weather conditions. From a network learning perspective, deep neural networks extract features at different levels through layer-by-layer feature extraction operations. Early layers capture low-level features such as edges, textures, and color gradients, while deeper layers learn more complex semantic information and scene structures. However, the loss of low-level features does not entirely compromise high-level scene information. The network can reconstruct the overall scene from remaining valid pixels or local features. Therefore, we argue that it is not necessary to perform image restoration first when handling complex scenes. Second, the task of multi-modality image fusion inherently involves discarding redundant information and retaining complementary data. Unlike image restoration, which aims to recover as much detail as possible without accounting for redundancy, fusion emphasizes retaining only the most significant information. By leveraging complementary multi-modality features, the network can concentrate on these useful details during reconstruction, rather than processing unnecessary pixel data. This not only reduces computational complexity but also enhances the ability of the network to recover the aspects of the scene that best represent its true characteristics. Thus, our objective is to shift the focus of the network from perfect pixel-level recovery to maximizing the representation of key scene information.
The fusion module was constructed to equip the network with the ability to extract multi-modality information, isolate the shared and unique features of each modality, and fuse them based on a unique prior for each feature type. Firstly, we introduce a learned low-rank representation (LLRR) block [4] to decompose the source images into sparse and low-rank components. This feature decoupling approach allows for a more comprehensive consideration of the feature information of the image. Additionally, the low-rank and sparse decompositions reduce the computational complexity of the fusion module. The low-rank component, with its fewer degrees of freedom, represents the global structure with fewer parameters, while the sparse component focuses on a limited number of salient pixels, minimizing unnecessary computational overhead. By leveraging this decomposition, more efficient algorithms can be applied to process large-scale images. For the features integration, a sparse feature prediction module based on an optimization model was proposed to reconstruct the sparse components, and a sparse transformer [20] was introduced to capture the long-range contextual information of an image.
In the restoration module, the clear feature prediction block was developed, guided by the atmospheric scattering model [21, 22]. This module derives illumination and reflectance features from images to predict the distribution of light transmission, utilizing learned convolutional sparse coding (LCSC) [23]. Thus, propelled by both mathematical optimization algorithms and physical models, our proposed model has interpretable and physically-aware properties. It obviates the necessity for model customization to cater to specific weather-related issues, and it enables the simultaneous training of multiple teacher networks under a unified framework for a singular weather restoration task. Subsequently, distillation learning was incorporated into the model to transfer knowledge from these diverse teacher networks to a singular student network, enabling the model to resolve degradation issues stemming from various severe weather phenomena.
@@
Furthermore, we introduce a large-scale benchmark, AWMM-100k, specifically designed for multi-modality image fusion in adverse weather conditions, aiming to advance research in this challenging field. Our dataset consists of 100,000 multi-modality image pairs, comprehensively covering both synthetic and real-world scenes affected by three major weather degradations: rain, haze, and snow. To ensure robust evaluation, the data spans multiple degradation levels for each weather type, capturing the diversity and complexity of natural environments. This benchmark addresses a critical gap in the field by providing a reliable, diverse, and extensive resource for developing and evaluating fusion methods under real-world adverse weather conditions.
The contributions of this study are summarized as follows.
@@
We propose an end-to-end all-weather MMIF framework, effectively fusing and recovering the perturbed multi-modality information. Unlike existing MMIF works, to the best of our knowledge, this is the first work to address MMIF and adverse weather removal simultaneously, defining a new challenging task.
@@
We propose a feature separation-based fusion module that considers the reconstruction of local and global features from sparse and low-rank components.
We propose a physically-aware clear feature prediction block (PFB) that comprehensively accounts for the restoration of clear features, thereby avoiding a time-consuming empirical network design.
Leveraging both existing MMIF datasets and our collected data, we built a dataset (AWMM-100k) dedicated to MMIF research under adverse weather. It containing 100,000 aligned infrared and visible image pairs with labels for object detection, establishing it as a versatile benchmark for a broad spectrum of vision tasks.
II Related Work
II-A DL-based MMIF Methods
@@
DL-based IVIF algorithms [24, 25, 26, 27, 28, 29, 30, 31] demonstrate robust fusion performance in conventional scenes, employing methodologies such as GAN or AE to effectively integrate salient feature information from each modality. With the widespread utilization of IVIF technology, researchers increasingly integrate this technology into downstream vision tasks [3, 32, 33, 34, 35]. Their aim is to enhance the extraction of multi-modality features, augmenting semantic information and leveraging fused images to optimize the performance of tasks such as object detection, semantic segmentation, and salient object detection. Furthermore, there are many joint multi-tasking algorithms designed to cope with real-world applications [36, 37, 38]. In practical applications, images captured by cameras may be slightly misaligned and distorted; some researchers have designed correction networks for fusion specifically for such situations [37]. Xu et al. [38] designed an IVIF model that could fuse unaligned image pairs, effectively reducing the cost of adopting IVIF technology. In some cases, the fixed input and output resolutions of images can limit the development of IVIF-based methods; Li et al. [39] designed an IVIF framework specifically for this problem. In addition, some algorithms apply the IVIF task to degraded environments, such as noise [40], low light [41], and low contrast scenes [42]. However, current research does not effectively address the generalization of IVIF algorithms, whether they are single-task or multi-task, to adverse weather conditions.
III Proposed Method
III-A Overview
As shown in the Fig. 1, AWFusion comprises two main parts: image fusion and image restoration. It is important to emphasize that the proposed algorithm is trained in an end-to-end manner. First, the multi-modality data affected by adverse weather is sent to the image fusion module for feature encoding. To effectively distinguish private and shared features from different modalities and thereby allow the fusion network to overcome the limitations in using a single encoder, a LLRR block is introduced. This block decomposes the source image into a low-rank component, which captures global structural information with high pixel correlation, and a sparse component, which represents high-frequency information with faster pixel variations. This approach enables the targeted handling of specific modal features and enhances fusion outcomes.
In the second part, the multi-modality interaction features obtained from the fusion module are fed into the restoration module to infer the original clear scene. As shown in Fig. 1, the restoration module in the model uses a U-Net backbone. In the backbone network, we stack five sparse transformer blocks (STB) to extract useful features distributed in adverse weather. STB mainly consist of top-k sparse attention (TKSA) and mixed-scale feed-forward network (MSFN). Regarding TKSA, it performs feature aggregation to keep the features in an image sparse, and MSFN extracts multiscale local information from the image [20]. The proposed PFB focuses on refining the spatial information, which is embedded into both sides of the backbone network and enforced before being inputted into subsequent modules, which helps recover the detailed information. Finally, the reconstructed fusion results are output through the restoration module.
Finally, drawing inspiration from Che et al. [43], we introduced a distillation learning approach in which distinct teacher models are initially trained for individual weather-related issues. Subsequently, knowledge is collaboratively transferred from these multiple teacher models to student models to comprehensively address weather-related challenges.
III-B Image Fusion Module
@@
Through the decomposition provided by the LLRR block, our network focuses solely on the fusion of low-rank and sparse features. Low-rank components reflect the global information and main structure of the image, constrained by certain rank features. Given the strong ability of Transformer to capture global context, we employ the STB to fuse low-rank features from different modalities. In contrast, sparse components represent local details such as edges, textures, and high-frequency information. To effectively manage these local fusions, we design the Sparse Feature Block (SFB).
III-B1 SFB
Sparse feature fusion is performed to effectively extract and integrate local features specific to an infrared image () and a visible image (). The fused feature map consists of two parts: the sparse feature for and the sparse feature for . is modeled as follows:
where and are the sparse filters of related to and , respectively, and denotes the number of filters. Subsequently, Eq. 1 can be transformed into the following optimization problem:
@@ III-B1 SFB
III-C Image Restoration Module
The second phase aims to achieve image restoration in the presence of three types of adverse weather conditions (rain, snow, and haze). Based on the atmospheric scattering model, adverse weather conditions such as rain, haze, and snow induce scattering and absorption of light by particles in the atmosphere. Consequently, these atmospheric phenomena lead to comparable image degradation across a spectrum of unfavorable conditions [44]. For instance, during heavy rainfall, the accumulation of distant rain lines and atmospheric water particles creates a ”rain curtain” effect. This visual effect of rainwater accumulation produces a haze like phenomenon in the image background [45]. The above conclusion provide a research basis to realize deweathering in the same framework. Therefore, for the restoration module, an atmospheric scattering model is combined to simulate the process of image degradation, and a PFB is proposed.
III-C1 PFB

### v3

@@ Abstract
Multi-modality image fusion (MMIF) combines complementary information from different image modalities to provide a comprehensive and objective interpretation of scenes. However, existing fusion methods cannot resist different weather interferences in real-world scenes, limiting their practical applicability. To bridge this gap, we propose an end-to-end, unified all-weather MMIF model. Rather than focusing solely on pixel-level recovery, our method emphasizes maximizing the representation of key scene information through joint feature fusion and restoration. Specifically, we first decompose images into low-rank and sparse components, enabling effective feature separation for enhanced multi-modality perception. During feature recovery, we introduce a physically-aware clear feature prediction module, inferring variations in light transmission via illumination and reflectance. Clear features generated by the network are used to enhance salient information representation. We also construct a large-scale MMIF dataset with 100,000 image pairs comprehensively across rain, haze, and snow conditions, as well as covering various degradation levels and diverse scenes. Experimental results in both real-world and synthetic scenes demonstrate that the proposed method excels in image fusion and downstream tasks such as object detection, semantic segmentation, and depth estimation. The source code is available at https://github.com/ixilai/AWFusion.
keywords:
@@
Image fusion, Adverse weather, Multi-modality benchmark, Image restoration1 Introduction
@@
Infrared and visible image fusion (IVIF) is performed to integrate the complementary and salient feature information from different modal images to achieve a more comprehensive interpretation of a target scene. By integrating data from different spectral ranges, IVIF contributes to advancements in computer vision tasks [32, 17, 38, 33, 28], including semantic segmentation [24, 3], object detection [54, 55], and unmanned driving [62].
Recently, deep learning (DL) has exhibited remarkable progress within the realm of artificial intelligence [65], owing to its exceptional aptitude for intricate nonlinear approximations and adeptness in extracting salient features. Consequently, several algorithms tailored to the IVIF have been developed. These algorithms can be broadly categorized into four groups: generative adversarial networks (GAN)-based models [18, 67], diffusion-based models [75, 69], autoencoder (AE)-based models [73, 53], and algorithmic expansion models [23, 10]. However, most of these methods are designed for ideal imaging conditions. With the increasing need for robust perception in real-world applications, IVIF is expected to deliver reliable scene representation under adverse weather, where visibility is reduced and modality information becomes severely imbalanced. Although current DL-based fusion methods [60, 2, 7, 30] achieve promising results in clear-weather scenarios, their generalization and robustness remain limited in rainy, hazy, or snowy environments. Therefore, three major challenges need to be addressed:
@@
Currently, there is a critical research gap in that IVIF cannot deliver scene information under adverse weather. Addressing this challenge requires more than isolated algorithmic solutions for specific weather-related issues, necessitating a unified framework to comprehensively restore intricate scene details affected by adverse weather.
@@
Generally, DL-based algorithms are driven by data. Existing datasets dedicated to collecting comprehensive imaging data from multiple modalities are complete in terms of scene diversity. However, no large datasets have been developed specifically for research on applying IVIF to scenarios with adverse weather conditions.
In summary, it is essential to build a large-scale multi-modality image dataset for specifically tailored for adverse weather scenarios, and to develop an effective unified “Restoration + Fusion” framework, enabling more reliable IVIF performance in challenging environments.
1.1 Our Contributions
To address above challenges, we propose an all-weather MMIF algorithm (AWFusion) that integrates both deweathering and fusion functions within a single framework, thereby providing an effective means to achieve robust fusion under adverse weather. To clearly present the design principles of the proposed algorithm, we describe it from the following three aspects.
@@ 1.1 Our Contributions
(1) Feasibility Analysis of Unified Framework: We argue that to construct a effective unified framework, complementary information should first be integrated without redundancy, followed by restoring the overall clear features. The main reasons are as follow: (1) In natural scenes, images contain sufficient redundancy, allowing the original content to be inferred even under adverse weather degradation. From a network learning perspective, deep neural networks extract features at different levels through layer-by-layer feature extraction operations. Early layers capture low-level cues like edges and textures, while deeper layers encode complex semantics and scene structures. However, the loss of low-level features does not entirely compromise high-level scene information. The network can reconstruct the overall scene from remaining valid pixels or local features. Therefore, it is not necessary to perform image restoration first when handling complex scenes. (2) The task of IVIF inherently involves discarding redundant information and retaining complementary data. Unlike image restoration, which aims to recover all details, image fusion focuses on preserving the most significant information. By leveraging complementary multi-modality features, the network emphasizes useful details while ignoring redundant pixels. This not only reduces computational complexity but also enhances the ability of the network to recover the aspects of the scene that best represent its true characteristics. Thus, our objective is to shift the focus of the network from perfect pixel-level recovery to maximizing the representation of key scene information. The unified framework is consisting of two main components: a fusion module and a restoration module.
(2) Implementation of Unified Framework: In the fusion module, we introduce a learned low-rank representation (LLRR) block [23] to decompose the source images into sparse and low-rank components firstly. This feature decoupling approach allows for a more comprehensive consideration of the feature information of the image. Additionally, low-rank and sparse decomposition reduces the computational cost of fusion module: the low-rank component can capture global structure with fewer parameters, while the sparse component focuses on salient pixels, enabling efficient processing of large-scale images. To integrate features, a Sparse Feature Prediction Block (SFB) was designed to reconstruct sparse components, and a Sparse Transformer Blocks (STB) [6] was introduced to capture the long-range contextual information of an image. In the restoration module, we design a the Physical-Aware Clear Feature Prediction Block (PFB), guided by the atmospheric scattering model [76]. The PFB derives illumination and reflectance features from images to predict the distribution of light transmission, utilizing learned convolutional sparse coding (LCSC) [46]. It removes the need for weather-specific model design, enabling simultaneous training of multiple teacher networks in a unified restoration framework. Subsequently, distillation learning transfers knowledge from multiple teacher networks to a single student network, enabling the model to handle adverse weather degradations.
@@
(3) Construction of Large-Scale All-Weather Benchmark: Furthermore, we introduce a large-scale benchmark, AWMM-100k, specifically designed for multi-modality image fusion in adverse weather conditions, aiming to advance research in this challenging field. Our dataset consists of 100,000 multi-modality image pairs, comprehensively covering both synthetic and real-world scenes affected by three major weather degradations: rain, haze, and snow. To ensure robust evaluation, the data spans multiple degradation levels for each weather type, capturing the diversity and complexity of natural environments. This benchmark addresses a critical gap in the field by providing a diverse, and extensive resource for developing and evaluating fusion methods under real-world adverse weather.
The significant contributions of this work are summarized below:
@@
We propose an end-to-end unified all-weather MMIF framework. To the best of our knowledge, this is the first work to address MMIF and adverse weather removal simultaneously, defining a new challenging task.
@@
We propose a feature separation-based fusion module that considers the reconstruction of local and global features from sparse and low-rank aspects, and develop a physically-aware feature prediction block that comprehensively accounts for the restoration of clear features.
We construct an all-weather MMIF benchmark (AWMM-100k) contains rain, haze, and snow scenes with different level. AWMM-100k covers real-world scenes, and has totally 100,000 registered multi-modality image pairs with object detection labels.
Extensive experiments on adverse weather image fusion, as well as downstream tasks cover depth estimation, semantic segmentation, and object detection, verify the effectiveness and superiority of the proposed method.
2 Related Work
2.1 Image Fusion in Conventional Conditions
DL-based IVIF algorithms [72, 44, 12, 13, 20, 36, 52, 27] demonstrate robust fusion performance in conventional scenes, employing methodologies such as GAN or AE to effectively integrate salient feature information from each modality. For example, Zhao et al. [74] introduced a self-supervised framework incorporating equivariant priors to address the lack of ground truth in multimodal image fusion. Similarly, Cheng et al. [7] exploited underlying visual tasks in digital photography fusion as supervisory signals and proposed a generalized fusion network that enhances cross-task interaction. With the widespread utilization of IVIF technology, researchers increasingly integrate this technology into downstream vision tasks [16, 25, 50, 39]. Their aim is to enhance the extraction of multi-modality features, augmenting semantic information and leveraging fused images to optimize the performance of tasks such as object detection, semantic segmentation, and salient object detection. For instance, Tang et al. [50] argued that most image fusion methods overlook the demands of high-level visual tasks, leading to a lack of semantic information in the fused results. To address this, they proposed a semantic-aware real-time image fusion network. Liu et al. [37] noted that existing approaches often treat fusion and downstream tasks as separate processes, which limits the benefits of fusion for subsequent tasks. To overcome this, they introduced a discriminative cross-dimensional evolutionary learning framework for IVIF, employing an evolutionary algorithm to optimize the loss function parameters.
2.2 Image Fusion in Complex Conditions
@@
Additionally, there are many joint multi-tasking algorithms designed to cope with real-world applications [21, 64, 56]. In practical applications, images captured by cameras may be slightly misaligned and distorted; some researchers [14] have designed correction networks for fusion specifically for such situations. Xu et al. [64] designed an IVIF model that could fuse unaligned image pairs, effectively reducing the cost of adopting IVIF technology. In some cases, fixed input and output resolutions constrain the development of IVIF methods. To address this limitation, Li et al. [19] To address this limitation, Li et al. proposed a meta-learning-based IVIF framework, where a meta-amplification module dynamically predicts amplification filter weights using scale factors as input. In addition, some algorithms apply the IVIF task to degraded environments, such as noise [29], low light [49], and low contrast scenes [68]. For example, Li et al. [31] proposed a unified IVIF framework for complex scenes based on optical-transmission decomposition. The framework leverages the prior that degradation factors cause scattering during light transmission, allowing the transmission map to capture the distribution of different degradations. By separating the degraded features from the source image, the loss of detail in the feature extraction process is avoided. Zhang et al. [70] combined multi-modal large models to develop a robust, language-driven image fusion algorithm. It leverages the denoising mechanism of latent diffusion models to progressively remove degraded information and can locate targets of interest based on language instructions, highlighting specific scene details. However, current research does not effectively address the generalization of IVIF algorithms, whether they are single-task or multi-task, to adverse weather conditions.
3 Proposed Method
3.1 Problem Formulation
Given an infrared and visible image pair captured under adverse weather, our objective is to generate a fused image that retains complementary modality information while recovering a clean visual scene. The task can be formulated as:
where denotes the proposed AWFusion network and represents the learnable parameters. This formulation involves two key goals: (1) effective cross-modal feature fusion and (2) restoration of weather-degraded structures and textures.
Fusion module. To disentangle complementary information, the input images are first decomposed into low-rank and sparse components using the LLRR block:
where denotes the processing performed by the Learned Low-Rank Representation Blocks, represent the global structure (low-rank components), and contain modality-specific salient details (sparse components).
The two components are then fused separately to better preserve shared structures and unique characteristics:
Where and denote the processing performed by the Sparse Transformer Blocks and the Sparse Feature Prediction Blocks, respectively, producing low-rank fused features and sparse fused features , which are then aggregated and fed into the restoration module for clear-scene reconstruction.
Restoration module. The restoration module employs a U-Net backbone augmented with five STB, each consisting of a Top- Sparse Attention (TKSA) mechanism and a Mixed-Scale Feedforward Network (MSFN) [6]. TKSA aggregates the most informative features while maintaining sparsity, and MSFN extracts multi-scale local information. Additionally, the PFB refines spatial information and is embedded on both sides of the backbone to improve feature representation. The restored fusion result is then output as the final prediction .
Finally, to enhance the robustness of model under adverse weather conditions, we introduce a distillation learning strategy [5]. Specifically, multiple teacher networks are first trained for distinct weather degradations, and their knowledge is then transferred to a single student network to comprehensively address adverse weather effects.
3.2 Image Fusion Module
@@
Due to significant differences between modalities and potential semantic conflicts in certain cases, directly fusing them in the spatial domain can lead to insufficient feature interaction, contrast loss caused by modality misalignment, or misalignment artifacts. Therefore, decomposing image features prior to fusion facilitates the network in capturing interactions among different types of features, enabling a better understanding of the semantic information. Therefore, we decompose the image into low-rank and sparse components through the LLRR blocks. Low-rank components reflect the global information and main structure of the image, constrained by certain rank features. Given the strong ability of Transformer to capture global context, we employ the STB to fuse low-rank features from different modalities. In contrast, sparse components represent local details such as edges, textures, and high-frequency information. To effectively manage these local fusions, we design the SFB.
3.2.1 Sparse Feature Prediction Block (SFB)
Sparse feature fusion is performed to effectively extract and integrate local features specific to infrared images () and visible images (). Here, we employ an optimization principle inspired by LCSC to guarantee that the learned sparse filters represent meaningful high-frequency image details. This is motivated by the physical observation that edges and textures dominate local image variations, and proper modeling is necessary to preserve them in the fused output.
The fused feature map consists of two parts: the sparse feature for and the sparse feature for , modeled as:
where and are the convolutional filters for and , respectively.
To derive a tractable optimization problem from the convolutional representation in Eq. 4, we first express the convolutions in matrix-vector form. Each filter and is converted into a Toeplitz matrix and , while the corresponding feature maps and are vectorized. This allows the fused feature to be equivalently written as a linear combination of sparse components:
Next, to ensure that the sparse features accurately reconstruct the original images, we introduce reconstruction fidelity terms for each modality, measured by the -norm of the difference between the source images and their reconstructions:
Finally, to encourage sparsity in the learned features and suppress redundant activations, we add an regularization term for each sparse feature. Combining the fidelity and sparsity terms naturally leads to a joint optimization problem that can be solved using convolutional sparse coding. The sparse feature responses are then obtained by solving:
@@ 3.2.1 Sparse Feature Prediction Block (SFB)
3.3 Image Restoration Module
The second phase aims to achieve image restoration in the presence of three types of adverse weather conditions (rain, snow, and haze). Based on the Atmospheric Scattering Model (ASM), adverse weather conditions such as rain, haze, and snow induce scattering and absorption of light by particles in the atmosphere. Consequently, these atmospheric phenomena lead to comparable image degradation across a spectrum of unfavorable conditions [77, 31]. For instance, during heavy rainfall, the accumulation of distant rain lines and atmospheric water particles creates a ”rain curtain” effect. This visual effect of rainwater accumulation produces a haze like phenomenon in the image background [66]. The above conclusion provide a research basis to realize deweathering in the same framework.
To leverage this unified representation, we construct the PFB guided by the ASM. Specifically, the network is designed to predict the key parameters of the ASM, including the atmospheric light and the transmission map, using learnable neural modules. By estimating these parameters, the network can effectively model the scattering and absorption effects induced by adverse weather. Subsequently, the clear image features are reconstructed by combining the predicted atmospheric light and transmission map, enabling the recovery of scene information in a physically meaningful and interpretable manner. This approach enables the network to learn from data while incorporating physical degradation priors.
3.3.1 Physical-Aware Clear Feature Prediction Block (PFB)